OpenStudy (anonymous):

Solve: 2(x + 8) − (x + 1) = 20

jimthompson5910 (jim_thompson5910):

2(x + 8) − (x + 1) = 20 2x + 16 − x - 1 = 20 x + 15 = 20 I'll let you finish

OpenStudy (anonymous):

Thanks

jimthompson5910 (jim_thompson5910):

yw